1.) Find the sequence of elementary matrices whose product is:
A=[1/2 -3
2 3/15] << 2 x 2 matrix


2.) Solve the given SLE using LU-Factorization:
x(sub 1) + x(sub 2) + x(sub 3) = 3
x(sub1) - x(sub 2) + 4x(sub 3) = 4
2x(sub 1) + 3x(sub 2) - 5x(sub 3) = 0
1
Expert's answer
2020-10-22T17:57:01-0400

(i) A="\begin{bmatrix}
 \frac{1}{2} & -3 \\
 2 & \frac{3}{15}
\end{bmatrix}"


Let I="\begin{bmatrix}
 1 & 0 \\
 0 & 1
\end{bmatrix}"

Applying IA=A

"\begin{bmatrix}
 1 & 0 \\
 0 & 1
\end{bmatrix}" "\begin{bmatrix}
 \frac{1}{2} & -3 \\
 2 & \frac{3}{15}
\end{bmatrix}" ="\begin{bmatrix}
 \frac{1}{2} & -3 \\
 2 & \frac{3}{15}
\end{bmatrix}"


Applying "R_1\to2R_1"

"\begin{bmatrix}
 2 & 0 \\
 0 & 1
\end{bmatrix}
\begin{bmatrix}
 \frac{1}{2} & -3 \\
 2 & \frac{3}{15}
\end{bmatrix}
=\begin{bmatrix}
 {1} & -6\\
 2 & \frac{3}{15}
\end{bmatrix}"


Applying "R_2\to R_2-2R_1"

"\begin{bmatrix}
 2 & 0 \\
 -2 & 1
\end{bmatrix}
\begin{bmatrix}
 \frac{1}{2} & -3 \\
 2 & \frac{3}{15}
\end{bmatrix}
=\begin{bmatrix}
 {1} & -6\\
 0 & \frac{183}{15}
\end{bmatrix}"


Applying "R_1\to R_1+\frac{90}{183}R_2"


"\begin{bmatrix}
 \frac{186}{183}& \frac{90}{183} \\
 -2 & 1
\end{bmatrix}
\begin{bmatrix}
 \frac{1}{2} & -3 \\
 2 & \frac{3}{15}
\end{bmatrix}
=\begin{bmatrix}
 {1} & 0\\
 0 & \frac{183}{15}
\end{bmatrix}"


applying "R_2\to \frac{15}{183}R_2"

"\begin{bmatrix}
 \frac{186}{183}& \frac{90}{183} \\
 \frac{-30}{183} & \frac{15}{183}
\end{bmatrix}
\begin{bmatrix}
 \frac{1}{2} & -3 \\
 2 & \frac{3}{15}
\end{bmatrix}
=\begin{bmatrix}
 {1} & 0\\
 0 & 1
\end{bmatrix}"



Sequence of elementary matrix are

"\begin{bmatrix}
 1 & 0 \\
 0 & 1
\end{bmatrix}" "\begin{bmatrix}
 2 & 0 \\
 0 & 1
\end{bmatrix}" "\begin{bmatrix}
 2 & 0 \\
 -2& 1
\end{bmatrix}\begin{bmatrix}
 \frac{186}{183} & \frac{90}{183} \\
 -2 & 1
\end{bmatrix}\begin{bmatrix}
 \frac{186}{183} & \frac{90}{183} \\
 \frac{-30}{183} & \frac{15}{183}
\end{bmatrix}" =A


(ii) Given equation are

"x_1+x_2+x_3=3"

"x_1-x_2+4x_3=4"

"2x_1+3x_2-5x_3=0"


A="\begin{bmatrix}
 1 & 1 &1\\
 1 & -1&4\\
 2 & 3 &-5
\end{bmatrix}" ,B="\begin{bmatrix}
 3 \\
 4\\
0
\end{bmatrix}" ,X="\begin{bmatrix}
 x_1 \\
 x_2\\
x_3
\end{bmatrix}"


Transforming matrix A to upper triangular matrix(U) using elementary row transformation and Lower triangular matrix(L) simultaneously

Applying "R_2\to R_2-R_1"


"U=\begin{bmatrix}
 1 & 1 &1\\
 0 & -2&3\\
 2 & 3 &-5
\end{bmatrix}" ,L="\begin{bmatrix}
 1 & 0 &0\\
 1 & 1&0\\
 0 & 0 &1
\end{bmatrix}"


Appling "R_3\to R_3-2R_1"

"U=\begin{bmatrix}
 1 & 1 &1\\
 0 & -2&3\\
 0 & 1 &-7
\end{bmatrix}" ,L="\begin{bmatrix}
 1 & 0 &0\\
 1 & 1&0\\
 2 & 0 &1
\end{bmatrix}"



Applying "R_3\to R_3+\frac{1}{2}R_2"


"U=\begin{bmatrix}
 1 & 1 &1\\
 0 & -2&3\\
 0 & 0 &\frac{-11}{2}
\end{bmatrix}" ,L="\begin{bmatrix}
 1 & 0 &0\\
 1 & 1&0\\
 2 & \frac{-1}{3}&1
\end{bmatrix}"


So LY=B

"\begin{bmatrix}
 1 & 0 &0\\
 1 & 1&0\\
 2 & 0 &1
\end{bmatrix}
\begin{bmatrix}
 y_1\\
 y_2\\
 y_3
\end{bmatrix}
=\begin{bmatrix}
 3\\
 4\\
 0
\end{bmatrix}"


"\begin{bmatrix}
 y_1\\
 y_1+y_2\\
 2y_1+y_3
\end{bmatrix}" ="\begin{bmatrix}
 3\\
 4\\
 0
\end{bmatrix}"


"y_1=2,y_1+y_2=4,y_2=2"


"2y_1+y_3=0, y_3=-2"


Now UX=Y


"\begin{bmatrix}
 1 & 1 &1\\
 0 & -2&3\\
 0 & 1 &-7
\end{bmatrix}
\begin{bmatrix}
 x_1\\
 x_2\\
 x_3
\end{bmatrix}
=\begin{bmatrix}
 2\\
 2\\
 -2
\end{bmatrix}"


"x_1+x_2+x_3=2, -2x_2+3x_3=2"

"x_2-7x_3=-2"


Solving these equation we get,

"x_1=\frac{28}{11},x_2=\frac{-8}{11},x_3=\frac{2}{11}"